OpenStudy (anonymous):

\[−(√3/2)^2+1=\cos2x\] cos x = 1/2 I'd like to know how the negative sign distributes to make it 1/2 and not -1/2. Can someone please explain?

OpenStudy (anonymous):

hey @amistre :)

OpenStudy (anonymous):

oh no it should = cos x not cos 2x

OpenStudy (amistre64):

\[-(a)^2=-(a*a) = -a*a\]hi

OpenStudy (amistre64):

1-3/4 = 1/4 cos(2x) = 1/4

OpenStudy (anonymous):

i typed it wrong

OpenStudy (anonymous):

−(√3/2)2+1=cos^2x cos x = 1/2 Sorry man!! that's how it should look

OpenStudy (anonymous):

−(√3/2)^2+1=cos^2x cos x = 1/2 <-- OMG i mean this one!!

OpenStudy (amistre64):

\[cos^2(x)=1/4\] sqrt each side cos(x) = 1/2

OpenStudy (anonymous):

ahh that makes sense, but what happens to the negative sign?

OpenStudy (amistre64):

its used

OpenStudy (anonymous):

it stays

OpenStudy (anonymous):

wont it as it is outside

OpenStudy (amistre64):

1 - (sqrt(3)/2)^2 = 1 - 3/4 = 1/4

OpenStudy (anonymous):

but it isnt 3/2 square it is root3/2 square

OpenStudy (anonymous):

see d question again

OpenStudy (anonymous):

if it's -(sqrt3/2)^2 + 1 = cos^2 x ----> how does it go to 1 - (sqrt(3)/2)^2 = 1 - 3/4 = 1/4 ? (sorry!)

OpenStudy (amistre64):

look up: commutative property of addition

OpenStudy (anonymous):

OK, so basically -(sqrt3/2)^2 + 1 = 1 - (sqrt3/2)^2 ?

OpenStudy (anonymous):

ya josh d ans comes out to be 1/4 for cos^2x but wen v root both sides the ans comes out to be 1/2

OpenStudy (amistre64):

\[a+b=b+a\] \[(-a)+b=b+(-a)=b-a\]

OpenStudy (anonymous):

d ans simple it comes out easily

OpenStudy (anonymous):

Thank you so much! I understand now